OpenStudy (anonymous):

I have 2 questions I cannot find the answer to, please help:

OpenStudy (anonymous):

\[^{11}\sqrt{x ^{5}}*\sqrt{x}\]

OpenStudy (anonymous):

and

OpenStudy (anonymous):

\[(x ^{4})^{3/5}\]

OpenStudy (anonymous):

Simplify. Express the product as a radical expression

jhonyy9 (jhonyy9):

these are very easy just you need using the property of exponents

OpenStudy (anonymous):

I would if I really knew and understood it

jhonyy9 (jhonyy9):

i think that you can calcule it easy sure

jhonyy9 (jhonyy9):

came on i like halp you understanding

OpenStudy (anonymous):

Ya, you helped, but I still dont know what to do.

jhonyy9 (jhonyy9):

you need to do what i have wrote before look when (x^a)^b =x^(a*b) so than in your case what you need to do ?

OpenStudy (anonymous):

ok so can you help me plug in the numbers?

jhonyy9 (jhonyy9):

see if x on exponent 1/2 mean squarroot x so than 11root(x^5) = x^(5/11) because how i have wrote to you before the denominator of a fractional exponent allways will be the indice of radical

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

but x^5/11 is not the finished answer, correct?

jhonyy9 (jhonyy9):

so than your first exercise will be x^(5/11) * x^(1/2) =x^(5/11 +1/2) = can you continue it ?

jhonyy9 (jhonyy9):

hey !!!

OpenStudy (anonymous):

I got 21/22..that is so not right. I have a feeling........ and hi?

jhonyy9 (jhonyy9):

so 5/11 +1/2 =((5*2) +(11*1))/22 =(10+11)/22 =21/22 so than x^(5/11) *x^(1/2 =x^(21/22) so what mean radical indice 22 from x on exponent 21

OpenStudy (anonymous):

I do not understand what to do on the last line you just posted.

jhonyy9 (jhonyy9):

this wann to be 22 radical(x^21) ok ?

OpenStudy (anonymous):

I do not understand

jhonyy9 (jhonyy9):

\[\sqrt[22]{x^(21)}\]

jhonyy9 (jhonyy9):

now ?

jhonyy9 (jhonyy9):

inside radical wann being x^(21)

OpenStudy (anonymous):

radial wann? I have not heard of that term...

jhonyy9 (jhonyy9):

so learn it radical indice 2 wann meaning squarroot

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

and cuberoot wann being radical indice 3

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so cuberoot(x)=radical indice 3 from x

jhonyy9 (jhonyy9):

so now do you understand your exercise

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):